# LoCo: Low-Contrast-Enhanced Contrastive Learning for Semi-Supervised Endoscopic Image Segmentation

### v1

@@ 4.3 Ablation Studies
Table 4 highlights several key findings regarding the performance of the model variants. M1 serves as the supervised baseline, utilizing only labeled data, and achieves a malignant IoU of 65.69%, a benign IoU of 36.65%, and a mIoU of 51.17%. M2 enhances these metrics by incorporating an additional loss function, $\mathcal{L}_{u}$, for unlabeled data. Building on M2, M3 introduces the proposed CDF module, leading to further improvements across all metrics. M4 and M5 enhance M2 by integrating BCE and ICE, respectively, resulting in noticeable gains. M6 combines both BCE and ICE with M2, yielding even better outcomes. Finally, M7 represents the complete LoCo model, which integrates the CDF module into M6, achieving the highest performance overall, with a malignant IoU of 72.19%, a benign IoU of 52.96%, and a mIoU of 62.57% among all variants.
4.4 Effectiveness of ICE
@@ 4.6 Effectiveness of CDF
4.7 Visualization Analysis under Various Label Partitions
@@
4.8 Segmentation Visualization Analysis

### v2

@@ IV-C Ablation Studies
Table IV highlights several key findings regarding the performance of the model variants. M1 serves as the supervised baseline, utilizing only labeled data, and achieves a malignant IoU of 65.69%, a benign IoU of 36.65%, and a mIoU of 51.17%. M2 enhances these metrics by incorporating an additional loss function, $\mathcal{L}_{u}$, for unlabeled data. Building on M2, M3 introduces the proposed CDF module, leading to further improvements across all metrics. M4 and M5 enhance M2 by integrating BCE and ICE, respectively, resulting in noticeable gains. M6 combines both BCE and ICE with M2, yielding even better outcomes. Finally, M7 represents the complete LoCo model, which integrates the CDF module into M6, achieving the highest performance overall, with a malignant IoU of 72.19%, a benign IoU of 52.96%, a mIoU of 62.57%, a DSC of 70.49%, and an NSD of 73.06% among all variants.
IV-D Effectiveness of ICE
@@ IV-F Effectiveness of CDF
IV-G Visualization Analysis under Various Label Partitions
@@
IV-H Segmentation Visualization Analysis
@@ V Conclusion
Acknowledgment
We would like to extend our heartfelt thanks to Yixuan Yuan and Yongcheng Li for their invaluable suggestions on enhancing the quality of this manuscript.